OpenStudy (anonymous):

If f(x) = 5√(2x), what is the value of f^-1(10)?

OpenStudy (anonymous):

\[f(x) = 5\sqrt{2x}. ------- f ^{-1}(10). Hope that makes \it easier \to visualize.\]

OpenStudy (anonymous):

you are trying to solve \[5\sqrt{x}=10\] for x

OpenStudy (anonymous):

oopx \[5\sqrt{2x}=10\] \[\sqrt{2x}=2\] \[2x=4\] \[x=2\]

OpenStudy (anonymous):

so since \[f(2)=10\] we know \[f^{-1}(10)=2\]

OpenStudy (anonymous):

So if \[f(x) = y\] then \[f ^{-1}(y) = x\]?

OpenStudy (anonymous):

Nvm got it

OpenStudy (anonymous):

x=2